OpenStudy (anonymous):

please help me!! find the area of this triangle if C= 30 degrees, b= 6in, and a= 7in

OpenStudy (campbell_st):

well it looks like you need \[A = \frac{1}{2} \times a \times b \times \sin(C)\] since C is the included angle. Just substitute and then evaluate. Hope it helps.

OpenStudy (anonymous):

that didnt get me the answer

OpenStudy (campbell_st):

wow... whats the answer...?

OpenStudy (anonymous):

i'm still not sure. theres 4 answers to pick from. . A) 21 in^2 B) 10.5in^2 C) 42in^2 and D) 21.7 in^2

OpenStudy (anonymous):

i may have done my math wrong but i didn't get any of those answers

OpenStudy (campbell_st):

well the answer I just calculated, using the formula I posted, is listed in you're choices...

OpenStudy (anonymous):

i can't get it. i'm an 11th grade math student so maybe that's why i'm not understanding it, and i don't have my teacher here to ask

OpenStudy (campbell_st):

so just draw option to show what your equation looks like when you substitute

OpenStudy (anonymous):

when i substituted everything in i got A= 1/2(7)(6)sin(30)

OpenStudy (campbell_st):

yep... thats correct... now use type it into a calculator, you can use 0.5 instead of 1/2 and see what you get

OpenStudy (anonymous):

21sin30

OpenStudy (campbell_st):

does your calculator have a button that allows you to convert to a decimal answer... like... opposite facing arrows or an S D button.... if so press it with the answer displayed

OpenStudy (anonymous):

i got it!! thank you sooo much. 10.5 right?

OpenStudy (campbell_st):

glad to help, yes and thats what I got

OpenStudy (anonymous):

would it be too much to help with another?

OpenStudy (campbell_st):

just post it and I'll look at it